### v1

@@ Abstract
Biological intelligence uses a “multiscale competency architecture” (MCA). It exhibits adaptive, goal directed behaviour at all scales, from cells to organs to organisms. In contrast, machine intelligence is only adaptive and goal directed at a high level. Learned policies are passively interpreted using abstractions (e.g. arithmetic) embodied in static interpreters (e.g. x86). Biological intelligence excels at causal learning. Machine intelligence does not. Previous work showed causal learning follows from weak policy optimisation, which is hindered by presupposed abstractions in silico. Here we formalise MCAs as nested “agentic abstraction layers”, to understand how they might learn causes. We show that weak policy optimisation at low levels enables weak policy optimisation at high. This facilitates what we call “multiscale causal learning” and high level goal directed behaviour. We argue that by engineering human abstractions in silico we disconnect high level goal directed behaviour from the low level goal directed behaviour that gave rise to it. This inhibits causal learning, and we speculate this is one reason why human recall would be accompanied by feeling, and in silico recall not.
@@
multiscale competency swarm intelligence consciousness.
@@ 1 Introduction
Biological intelligence is sample efficient. It is able to generalise or infer cause after only a few examples. Causal learning facilitates theory of mind, subjective experience and self awareness [bennett2023c, barron2016]. Our purpose here is to better understand how this capability might be engineered. Previous work established that causal learning requires one construct rather than presuppose variables and objects [bennett2023c]. By optimising for policies that express “weak” constraints on behaviour, relations between “contentless” [thompson2016] states are reduced to objects which cause affect [bennett2023b]. These claims were supported by proofs and experiments, but the resulting system is inefficient because it relies upon a high level, static abstraction layer. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. The abstraction layer on which the aforementioned experiments were run conveys an inductive bias. It limits what sort of “causes” can efficiently be learned. How might we achieve efficient causal learning? What sort of abstraction layer would fix this problem?
1.0.1 Biological multiscale competency architecture:
@@
Biology exhibits the capability we desire, so we look to biology for a more suitable abstraction layer. In the 70 years since psychoneural identity theory was proposed [sep-mind-identity], it has become evident that biological intelligence is not neatly confined to a neuronal abstraction layer [ciaunica2023]. Cells are now understood to be an agentic material [ball2023] and biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [morokuma2017], tadpoles with eyes on their backs that see [blackiston2013], and biological “xenobots” that perform simple tasks [levin2020]. Biology uses a “multiscale competency architecture” (MCA), in which lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [fields2020]. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Cancer is an example of what happens when this fails [levin2021], as cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [davies2011]. MCA is relevant to AI because biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [reid2011]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, human intelligence is the high level goal directed behaviour of a collective of cells. If we want to engineer more human-like machine intelligence, then we should mimic an MCA.
1.0.2 Agentic Abstraction Layers:
High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. emphasised the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [ciaunica2023b]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above, like how a CPU based on the x86 instruction set architecture (ISA) would the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts, compelling the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [thompson2007] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate higher level goals.
1.0.3 Summary:
Do MCAs facilitate causal learning? How? Previous AGI research:
@@ 1.0.3 Summary:
    formalised enactive cognition, to describe goal directed behaviour in terms about which objective claims can be made [bennett2023b, bennett2023a],
@@
    proved that a weaker policies are optimal in terms of sample efficiency [^1] Also demonstrating a $110-500\%$ improvement in causal learning in experiments. [bennett2023b], causal learning, and
@@
    argued this facilitates theory of mind, reafference and self awareness [bennett2023c].
Here we:
  1. 1.
    examine the behaviour of MCAs in the context of this formalism, and show weaker policies at higher levels imply weaker policies at lower levels, facilitating what we call multiscale causal learning.
  2. 2.
    We argue a static ISA like x86 hinders causal learning in comparison, and
  3. 3.
    that lower levels “compel” higher. We speculate this may explain why human memory and feeling are intertwined. By disconnecting high level from low in silico, we make it possible to interpret without the impetus of “feelings”.
2 The Formalism
The following definitions are shared with [bennett2024b, bennett2023a]. More extensive explanations of these definitions are given there. We have endeavoured to make key aspects of this paper understandable without the formalism, so readers who wish to should simply skip to the next section. The definitions are here for those who wish to dig into the details.
Definition 1 (environment)
@@ 2 The Formalism
Definition 2 (abstraction layer)
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless explicitly stated otherwise. If $\mathfrak{v}=P$, then we say that there is no abstraction.
@@
(notation) $E$ with a subscript is the extension of the subscript222e.g. $E_{l}$ is the extension of $l$..
@@
Definition 3 ($\mathfrak{v}$-task)
For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
@@
  • An output is a completion of an input [def. 2].
@@ 2.0.1 Learning and inference definitions:
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [bennett2023b]).
@@
Definition 6 (uninstantiated-tasks)
The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7 (utility of intelligence)
Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
@@
3 Arguments
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed, and how “the behaviour of subunits percolates up” to higher level goal directed behaviours, which in turn “constrains and facilitates” lower level behaviours [levin2024]. Once that is done we show how an MCA can facilitate more efficient causal learning than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
@@ 3 Arguments
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$. The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [bennett2023c] that differentiates observation from intervention in the manner of reafference [barron2016].
@@
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy. This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [bennett2023b, prop. 1,2]. For our cell example, weak policy at the level individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words the construction of a weak policy is not just enabled by an MCA, but is required to form one at all. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; an emergent property we ascribe to the behaviour of the lowest level. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [bennett2023b, proposition 1,2] facilitates causal learning at a single scale [bennett2023c], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
@@
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 7) for tasks at higher levels, and thus the efficiency of causal learning. That a collective of cells is an “organ” is just an abstract classification we ascribe to the collective’s behaviour. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, an “abstraction layer” is a property we ascribe to the emergent behaviour of many parts interacting. It appears to us as a second order system necessarily implied by the outputs $\mathfrak{v}$-tasks at different scales, because we see the whole rather than the parts. In the above example, $\omega$ is just $\alpha_{1}$ and $\alpha_{2}$ together, but to understand how we might “intervene” in the causal chain and disconnect emergent properties from the system which gave rise to them, we need to abstract away the individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$. We use this fact to justify representing policies abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
@@
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour which has emergent properties. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Proposition 1 (propagation)
@@
When we say we want efficient causal learning in computers, what we’re saying is that we want to mimic the adaptability of high level goal directed behaviour in humans. Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
@@
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.

### v2

@@ Abstract
Biological intelligence is more sample-efficient than artificial intelligence (AI), learning from fewer examples. Here we answer why. Given data, there can be many policies which seem “correct” because they perfectly fit the data. However, only one correct policy could have actually caused the data. Sample-efficiency requires a means of discerning which. Previous work showed sample efficiency is maximised by weak-policy-optimisation (WPO); preferring policies that more weakly constrain what is considered to be correct, given finite resources. Biology’s sample-efficiency demonstrates it is better at WPO. To understand how, we formalise the “multiscale-competency-architecture” (MCA) observed in biological systems, as a sequence of nested “agentic-abstraction-layers”. We show that WPO at low levels enables synthesis of weaker policies at high. We call this “multiscale-causal-learning”, and argue this is how we might construct more scale-able, sample-efficient and reliable AI. Furthermore, a sufficiently weak policy at low levels is a precondition of collective policy at higher levels. The higher level “identity” of the collective is lost if lower levels use an insufficiently weak policy (e.g. cells may become isolated from the collective informational structure and revert to primitive behaviour). This has implications for biology, machine learning, AI-safety, and philosophy.
@@
multiscale competency, AI, causal learning, sample efficiency
@@ 1 Introduction
Intelligence is often understood as predicting the outputs of a system, given inputs [1, 2, 3]. In that frame, learning is the act of leveraging data [^1] A set of inputs and outputs known to correspond to a particular system. We also refer to this as training or historical data. to improve predictive accuracy. It is an optimisation problem. The cause of data is the policy [^2] By policy, we just mean a constraint on outputs and inputs. In other words, a policy is a declarative definition of behaviour. There is an equivalence between imperative and declarative programs [4] so intuitively, consider a policy to be a program that dictates behaviour. We use the word policy for the sake of consistency with related work and to avoid having to use two different terms for the same thing. We do not mean only the more narrow definition of policy used in reinforcement learning. which originally generated it. To borrow a philosophical term [5], it is the policy which was the original “causal intermediary” between outputs and inputs. In machine learning, training data are sets of examples of goal directed behaviour; an ostensive definition of how to complete a task. A learning algorithm synthesises a policy which “fits” training data. Intuitively, any policy that fits is “correct” for that training data. However, the fact that a policy is correct with respect to data is not enough information to determine whether that policy was the cause of that data. Data could be the result of any number of policies, and if one wishes to accurately predict outputs given inputs that aren’t in the training data, then one must identify that policy which originally generated the data. The true cause. The closer one’s policy is to the cause, the more accurately one will be able to predict outputs given new inputs one hasn’t seen before. Conversely, the better one is at identifying the likely causes of data, the less data one will require to learn how to complete the task. The less data one requires, the more “sample efficiently” one learns.
1.1 Active Inference is Not Enough
Active inference [6, 7] formalises how organisms learn and adapt. It describes living systems in terms of a machine learning approach called variational Bayes. Like other machine learning approaches, this treats learning as a process which synthesises a policy which happens fit data. A policy that is correct relative to historical data, but may not be the true cause of that data. Yet biological intelligence (BI) is far more sample efficient than artificial intelligence (AI). Organoids learn far quicker than deep reinforcement learning agents in silico [8]. Yes, a robust AI will eventually learn cause and effect [9], but for now biology remains for more efficient. How is it that BI is so much more sample efficient than AI? This suggests there is more to the picture than active inference. BI is not just synthesising a policy which “fits” historical data; it is going a step further to figure out which among correct policies are likely to have caused historical data. How?
1.2 Weak-Policy-Optimisation is Causal Learning
Previous work already demonstrates how optimal sample efficiency [10] and causal learning [11] are achievable. Every policy implies a constraint on possible outputs and inputs; a constraint on how many possible tasks [^3] A task means a set of inputs and a corresponding set of outputs. A task is an ostensive definition of a goal or purpose. This is formally defined later in the paper. will be completed by this policy. A more specific policy completes fewer possible tasks, while a policy that implies a weaker constraint completes more possible tasks [12]. Past research into causal learning and sample efficiency found that the correct policies most likely to have caused historical data imply the weakest [^4] Formal definitions for policy weakness are given in the next section. constraint [10, 11] [^5] See Bennett’s Razor and related proofs in [10] for sample efficiency, and [11] for causality arguments.. Experiments demonstrated that a preference for weak policies correctly identifies the cause of historical data at $110-500\%$ the rate of a complexity based induction system [10]. Gradient descent does not optimise for weak policies, which is one reason why deep learning is so sample inefficient [13]. These results are further supported by the contravariance principle proposed by Cao and Yamins [14]. They argue tasks that more tightly constrain the space of correct policies [^6] A larger, more varied set of inputs and outputs implies a smaller set of correct policies. make it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. Given the sample efficiency exhibited by organoids [8], it would appear biology is far better at optimising for weak policies than current approaches to machine learning. Causal learning via weak-policy-optimisation [^7] Meaning learning using the weakness proxy as formally defined in the next section and in [3] (WPO) is clearly possible in silico, but existing implementations [10] are computationally expensive. One goal of our research is to better understand how the efficiency we observe in biology might be engineered.
1.2.1 How Weak Policy Optimisation Constructs Causal Identities
Previous work also established that causal learning requires one construct rather than presuppose variables and objects [11]. Causal identities are learned by constructing $X$ such that $X$ causes $Y$, not by presupposing the environment is divided into objects including $X$ and then intervening to test if $X$ causes $Y$ [15]. By optimising for policies that express “weak” constraints on behaviour, otherwise meaningless primitives888e.g. Declarative programs the truth value of which can be known, but for the truth conditions can never be known because environmental states are observable only partially and indirectly [3, 16, 15]. or very low level abstractions may be clustered together to form “causal identities” for higher level abstractions, like variables [11]. The “do” operator which is used to distinguish intervention from observation in causal inference problems can be substituted by introducing variables for interventions [17]. That is why one can learn cause and effect via WPO [11], by constructing causal identities for interventions.
1.2.2 How Causal Identities Relate to Abstraction Layers
The construction of causal identities is a process by which an environment can be broken up into abstract classifications that are inherently meaningful (motivated and relevant to a goal). In philosophical terms, if one is motivated by attraction and repulsion from “contentless” [18] states [^9] Contentless meaning categorised, value-less, monolithic states with no discernible internal structure., then WPO is a means by which those states can be reduced to objects or events which cause the affect which motivates an organism [15]. In computational terms, this amounts to constructing an abstraction layer that is most fit for a purpose [3, 16]. For example, the x86 instruction set architecture (ISA) has hard coded interpretations of certain sorts, such as numbers and arithmetic operations [^10] These amount to policies by the formal definition we use [3, 16]. As a result, the ISA can more efficiently address some problems than others. It is for this reason field programmable gate arrays exist as an alternative to static ISAs like x86; so that specific architectures can be constructed to more efficiently address specific problems.
1.2.3 Static Abstraction Layers and Computational Dualism
The ability to construct causal identities is easily hindered by the presupposition of static abstractions [11, 3]. This is because one cannot construct weaker abstractions if one is forced to use a static set of high level abstractions, like numbers and arithmetic operations. This is symptomatic of problem called “computational dualism[3]. Where 16th century Cartesian interactionist dualism proposed separate mental and physical substances interacting through the pineal gland, computational dualism is the concept of software distinct from hardware. For example, Hinton described software as “immortal”, existing infinitely many times across devices [19]. According to the argument against computational dualism [3], immortal computations do not exist. There are only finitely many instances of mortal computations. Intuitively, when one loads software onto a computer, no object moves into the computer. All that happens is that the state of hardware changes. Addressing computational dualism revealed weakness to be the confounding variable that correlates simplicity with generalisation [16, Prop. 2], allowing objective claims to be made regarding theoretical intelligent agents [3], addressing the subjectivity problems of complexity based induction systems in interactive settings [20]. This is all relevant to us because static abstraction layers are based on computational dualism. They seem to suggest software is something other than the state of hardware. Something interchangeable. Move-able. This means static abstraction layers are necessarily one size fits all, rather than optimised for particular purpose. The inefficiency introduced at the abstraction layer affects everything built within it, including any software based machine learning systems. Intuitively, an abstraction layer is like a window through which only part of the world can be seen. It conveys an inductive bias. Presupposing the world is formed of numbers, objects and operations of a specific sort limits which cause and effect relations can efficiently be learned. The sample efficiency exhibited by biology suggests it uses a more versatile alternative to static abstraction layers based on computation dualism; one more suitable for WPO.
1.3 Multiscale Competency Architecture
@@
Biology is slowly becoming a matter of engineering. It is now possible to create worms with two heads [21], tadpoles with eyes on their backs that see [22], and biological “xenobots” that perform simple tasks [23]. As biology becomes more tractable, the similarities between biology and human artifice (e.g. computers) have become more obvious. Cells are now understood to be an agentic material [24] for the purpose of engineering. They are but one scale in the “multiscale competency architecture” (MCA) of biology. In an MCA lower level goal directed behaviours determine higher level behaviours, and higher level behaviours constrain lower [25]. Adaptation takes place at every scale. For example, cells have a “perceptual field”, and can co-operate with other cells in their perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [26]. Cancer happens when this fails [27]. Cells isolated from the informational structure of the collective lose their high level identity, and pursue lower level goals [28]. Biological collectives exhibit high level goal directed behaviours far beyond the capability of any individual part [29]. Cells form organisms which in turn form yet more collectives up to the level of an entire ecosystem. In other words, there is a great deal of evidence [26] to suggest that human intelligence is nothing more than the high level goal directed behaviour of a collective of cells. If we want sample-efficient AI, then we should mimic an MCA.
2 Materials and Methods
To summarise what has been covered so far:
@@ 2 Materials and Methods
    Our goal is to understand why BI is so much more sample efficient (able not just to synthesise a policy that “fits” data, but to infer which among such correct seeming policies originally generated or caused the data).
@@
    As sample efficiency is maximised by WPO, our purpose is to understand how biology is able to synthesise weak policies so effectively.
@@
    A promising avenue appears to be the MCA [25] used by biological systems, instead of the static abstraction layers used in conventional computers (which exhibit the shortcomings of computational dualism [3]).
Hypothesis: The agentic, nested structure of MCAs allows more effective and efficient WPO than would be possible within a static layer like x86.
2.1 Agentic Abstraction Layers
To begin to understand how this might be possible, we must formalise MCAs in rigorous, mathematical terms. High level goal directed behaviour can emerge from distributed systems like swarms of cells, but it is extremely difficult to precisely control this sort of emergent high level goal directed behaviour. Work by Ciaunica et. al. stressed the importance of nested collectives in biology, exploring the relation between mother and child (each being a collective with their own needs and wants, and a collective together) during pregnancy [30]. Taking our cue from that approach, we propose formalising MCAs as nested and agentic abstraction layers. These abstraction layers are descriptions of behaviour, so we need not concern ourselves with the exact details of the system. It could be like biology, or something else. Our primary concern is how we might engineer and control efficient causal learning. A non-agentic, static abstraction layer passively interprets the abstraction layers above; like how a CPU based on the x86 instruction set architecture (ISA) interprets the behaviour of a python interpreter. By an agentic abstraction layer, we mean one that has its own agenda. It adapts. It compels the layers above to serve its own ends. Higher levels are compelled by lower to learn a policy “aligned” with theirs. Lower levels change their interpretation to enact [31] policy through higher levels, and are constrained by the resulting higher level behaviour. Information passes in both directions. Intuitively, if an abstraction layer is like a window through which only part of the environment can be seen, then an agentic abstraction layer is a window that moves on its own, to find the ideal view to facilitate the pursuit of higher level goals.
2.2 Definitions
We begin with a formalism [3] of enactive cognition [31]. We will use it to formalise MCAs in the next section. More extensive explanations of these definitions are given elsewhere [16, 3], and we have endeavoured to make key aspects of this paper understandable without the formalism; readers who wish to should simply skip to the next section.
Definition 1. [environment].
@@ 2.2 Definitions
Definition 2. [abstraction layer] .
  • We single out a subset $\mathfrak{v}\subseteq P$ which we call the vocabulary of an abstraction layer. The vocabulary is finite unless stated otherwise [^11] This is because a bounded system can contain only a finite amount of information [32] in a spatially extended environment like our own [16].. If $\mathfrak{v}=P$, then we say that there is no abstraction.
@@
(notation) $E$ with a subscript is the extension of the subscript121212e.g. $E_{l}$ is the extension of $l$..
@@
Definition 3. [$\mathfrak{v}$-task] For a chosen $\mathfrak{v}$, a task $\alpha$ is a pair $\langle{I}_{\alpha},{O}_{\alpha}\rangle$ where:
@@
  • An output is a completion of an input [def. 2.2].
@@ 2.2.1 Learning and inference definitions:
(optimal proxy) There is no proxy more sample efficient than $<_{w}$, so we call $<_{w}$ optimal. This formalises the idea that “explanations should be no more specific than necessary” (see Bennett’s razor in [10]).
@@
Definition 6. [uninstantiated-tasks] The set of all tasks with no abstraction (meaning $\mathfrak{v}=P$) is $\Gamma_{P}$ (it contains every task in every vocabulary). For every $P$-task $\rho\in\Gamma_{P}$ there exists a function $\lambda_{\rho}:2^{P}\rightarrow\Gamma_{P}$ that takes a vocabulary $\mathfrak{v}^{\prime}\in 2^{P}$ and returns a highest level child $\omega\sqsubset\rho$ which is also a $\mathfrak{v}^{\prime}$-task. We call $\lambda_{\rho}$ an uninstantiated-task, and $\lambda_{1}\sqsubset\lambda_{2}$ iff $\lambda_{1}(P)\sqsubset\lambda_{2}(P)$.
Definition 7. [utility of intelligence] Utility is the difference in weakness between the weakest and strongest correct policies of a task. The utility of a $\mathfrak{v}$-task $\gamma$ is $\epsilon(\gamma)=\underset{{\pi}\in{\Pi}_{\gamma}}{\max}\left(\lvert E_{\pi}\rvert-\lvert{O}_{\gamma}\rvert\right)$.
@@
3 Arguments and Results
We must first explain how $\mathfrak{v}$-tasks can represent MCAs. We must explain how different layers of the MCA are formed; how “the behaviour of subunits percolates up” to determine higher level behaviour, which in turn “constrains and facilitates” lower level behaviours [26]. Once that is done we show how an MCA can facilitate more efficient WPO than a static ISA. We conceive of an MCA as a causal ‘chain’ where each node is the goal directed behaviour of the node that preceded it. For example, cells form organs which form organisms.
@@ 3 Arguments and Results
We make no distinction between an object and its behaviour. Embodied and embedded, behaviour is limited to a particular vocabulary $\mathfrak{v}$ (meaning there the number of possible configurations is finite). The behaviour of two cells as individuals is two $\mathfrak{v}$-tasks $\alpha_{1}$ and $\alpha_{2}$. The collective behaviour of the two cells is a third $\mathfrak{v}$-task $\omega$ s.t. $\alpha_{1}\sqsubset\omega$ and $\alpha_{2}\sqsubset\omega$. Intuitively, this just means the two cells are parts of a larger whole. If there exists a correct policy for $\omega$, then the cells can behave in service of a common goal and thus have a common identity (for example if both cells use policies in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$). Note this identity is a precursor to the “causal identity for self” [11] that differentiates observation from intervention in the manner of reafference [33].
@@
If the parts of a larger whole learn, then the probability that they learn a shared policy is maximised if they use the weakness proxy (what we have been referring to as WPO). This is because weaker policies are more likely to generalise between $\mathfrak{v}$-tasks [10, prop. 1,2]. For our cell example, a sufficiently weak policy at the level of individual cells is necessary for the collective to have an identity, because only the weakest policies in $\Pi_{\alpha_{1}}$ and $\Pi_{\alpha_{2}}$ will also be in $\Pi_{\omega}=\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$. In other words WPO is not just facilitated by an MCA, but required for high level behaviour. To explain how higher levels constrain lower, consider that higher levels in an MCA are just an abstraction; something we ascribe to the behaviour of the lowest level. For the sake of intuition; different levels are patterns within one larger pattern, and we simplify the larger pattern by dividing it up. For example, a collective of cells remains cells whether we call it an organ or not. A collective does not exist in the absence of its parts. $\omega$ and $\alpha_{1}$ and $\alpha_{2}$ all exist in the same vocabulary. If the outputs of $\alpha_{1}$ affect the inputs of $\alpha_{2}$ and vice versa, then the outputs of $\omega$ determine the inputs of $\alpha_{1}$ and $\alpha_{2}$. The collective constrains behaviour at the lower level. We point this out to show how a $\mathfrak{v}$-task can represent an MCA. Just as the weakness proxy [10, proposition 1,2] facilitates causal learning at a single scale [11], in the context of an MCA it will factilitate causal learning at multiple scales, because the MCA as a whole is just another $\mathfrak{v}$-task.
@@
Intervening in the causal chain: We now illustrate how it is possible to “intervene” in the causal chain to disconnect high level behaviour from low level. We show this reduces “utility” (def. 2.2.1) for tasks at higher levels, meaning reduced capacity for WPO. We’re interested in understanding how MCAs relate to abstraction layers in a computer. Just as an organ is really just cells, x86 is a classification we ascribe to the product of human behaviour. It is a higher level behaviour of an MCA that includes humans. To understand how we might “intervene” in the causal chain and disconnect higher level behaviours from the system which gave rise to them, we need to abstract away the lower levels. To relate this back to the cell example, we must abstract away individual cells and focus on the consequences of their behaviour. A policy $\pi\in\Pi_{\alpha_{1}}\cap\Pi_{\alpha_{2}}$ implies an abstraction layer in the sense that for every $o\in E_{\pi}$ there exists $f\in P$ s.t. $\bigcap o=f$ (we can substitute $o$ with a declarative program $f$ that has the same extension and thus truth conditions). We use this fact to justify representing policies as abstraction layers, using a function $\mathfrak{f}$ that takes the extension $E_{\pi}$ of a collective’s policy $\pi$, and returns a new vocabulary $\mathfrak{v^{\prime}}=\{f\in P:\exists o\in E_{\pi}(\bigcap o=f)\}$. We can then use a sequence of uninstantiated tasks $\langle\lambda^{0},\lambda^{1}...\lambda^{n}\rangle$ s.t $\lambda^{i+1}\sqsubset\lambda^{i}$ to represent selection pressures that define “correctness” at every level. At each successive level, “correct” gets more specific. Because everything that happens in a higher level is correct according to lower levels, each higher level must further constrain “correct” if it is to exist in any meaningful sense.
@@
To construct layers we would start with a vocabulary $\mathfrak{v}^{i}$ at $i=0$, for example implied by a collective of cells or a lower level system. If this system exhibits goal directed behaviour, then it is given by the $\mathfrak{v}^{i}$-task $\lambda^{i}(\mathfrak{v}^{i})$ and a policy $\pi^{i}\in\Pi_{\lambda(\mathfrak{v}^{i})}$. $\pi^{i}$ implies behaviour. We translate policy into an abstraction layer using the aforementioned $\mathfrak{f}$. $\mathfrak{f}$ is applied to $E_{\pi^{i}}$ to obtain a vocabulary $\mathfrak{v}^{i+1}$, and then $\lambda^{i+1}$ is applied to $\mathfrak{v}^{i+1}$ to obtain a task $\lambda^{i+1}(\mathfrak{v^{i+1}})$. If $\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}\neq\emptyset$, then there exists a common purpose or “self” at this level of abstraction and we select a another policy $\pi^{i+1}\in\Pi_{\lambda^{i+1}(\mathfrak{v^{i+1}})}$.
Theorem 1.
@@
Humans are not just part of MCAs. A human is an MCA. Human abstractions are emergent, behavioural properties of that MCA.
@@
When we create an abstraction layer like the ISA of a computer, we create a machine that mimics our more abstract behaviours. However, the ISA is fixed. It does not learn or adapt the way the lower levels of the human MCA does. It has no agency, the way cells do, because it is not adapting in pursuit of a goal. No graceful degradation. Once we have built an ISA, it exists in the environment independently of the human MCA from which it emerged. This interrupts the causal chain.